the telephone rates between your city and the city where your brother lives are $.54 for the first minute and $.22 for each additional minute. calculate the cost of a 10-inute call to your brother.
Answer by LilSkittleMd(119)

0.54+9(0.22)=
0.54+1.98=2.52
The cost of a 10 minute phone call to your brother would be $2.52